PRENTER VERSION BACK Question 19 The Zin Company wants to achieve a profit of $126,000. They currently sell 5,000 units at a price of $72 per unit with variable costs of $32 per unit; fixed costs total $84,000. In order to achieve their target profit without increasing the selling price, the company will need to O increase sales by 250 units O increase sales by 600 units keep sales at their current levels O increase sales by $84,000 LINK TO TEXT Ouestinn
Answer #1

Target profit = $126,000

Selling price = $72 per unit

Variable cost = $32 per unit

Fixed costs = $84,000

Current sale = 5,000 units

Contribution margin per unit = Selling price per unit - Variable cost per unit

= 72 - 32

= $40

Number of units to be sold to earn the target profit = (Fixed costs + Target profit)/Contribution margin per unit

= (84,000 + 126,000)/40

= 210,000/40

= 5,250

Hence, sales must be increased by 250 units (5,250 - 5,000) to achieve the target profit

Correct option is (a)

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubts. Thanks.